OpenStudy (lgbasallote):

show that \[\mathcal L \lbrace \cos kt \rbrace = \frac{s}{s^2+k^2}; \; \text{for} \; s >0\] what does this mean?

OpenStudy (anonymous):

You have to show that \[ \int_0^\infty e^{- s t} \cos( k t)dt=\frac {s}{s^2 + k^2} \]

OpenStudy (lgbasallote):

what does s>0 mean?

OpenStudy (anonymous):

Try to write \[ \cos(k t) =\frac 1 2 (e^{i k t} + e^{-i kt}) \]

OpenStudy (lgbasallote):

where did that come from?

OpenStudy (anonymous):

De Moivres Formula

OpenStudy (lgbasallote):

hmm...we werent taught that...i think

OpenStudy (turingtest):

The Laplace transform of a function \(f(x)\) is defined as\[\mathcal L\{f(t)\}=\int_0^\infty e^{-st}f(t)dt~;~~~~~s>0\]if we did not put the restraint that \(s>0\) our integral would usually not converge (it would be line ignoring the negative sign)

OpenStudy (anonymous):

or you can use http://www.wolframalpha.com/input/?i=integrate+e^%28+a+x%29++cos%28+b+x%29

OpenStudy (turingtest):

since all you asked was what it meant that's all I answered @eliassaab has shown you how to find your particular transform by using the definition

OpenStudy (turingtest):

I usually resort to using a Laplace table....

OpenStudy (anonymous):

May be they want to see how to get this value in the table.

OpenStudy (lgbasallote):

oh so s>0 doesnt really matter much in the solutionn?

OpenStudy (anonymous):

If s <0, the integral will not converge.

OpenStudy (lgbasallote):

yeah i meant it's just a condition to ensure that it converges

OpenStudy (lgbasallote):

so how will i prove this? withoout the use of that de moivres thingy...coz we werent taught that

OpenStudy (anonymous):

You have to do integration by parts twice.

OpenStudy (lgbasallote):

ok i'll try

OpenStudy (anonymous):

http://www.youtube.com/watch?v=KTerp411MuM

OpenStudy (lgbasallote):

what is \[\int e^{-st}?\]is it just \[-\frac{e^{-st}}{s}\]

OpenStudy (anonymous):

Yes

OpenStudy (lgbasallote):

thanks..wanted to make sure s was a onstant

OpenStudy (lgbasallote):

constant*

OpenStudy (anonymous):

yes it is a constant.

OpenStudy (lgbasallote):

so i got to \[\large \int e^{-st} \cos kt dt + \frac{1}{s^2} \int e^{-st} \cos kt dt= -\frac{e^{-st} \cos kt}{s} +\frac{e^{-st} \sin kt - 1}{s^2} \] so how do i combine that?

OpenStudy (turingtest):

you didn't do it quite right

OpenStudy (lgbasallote):

aww

OpenStudy (turingtest):

\[\int e^{-st}\cos(kt)dt\]\[u=e^{-st}\implies du=-se^{-st}dt\]\[dv=\cos(kt)dt\implies v=\frac1k\sin(kt)\]so\[\int udv=uv-\int vdu\]\[=\frac1ke^{-st}\sin(kt)+\frac sk\int e^{-st}\sin(kt)dt\]now just do the second integral...

OpenStudy (lgbasallote):

wait...shouldnt you use trig functions instead of exponential as u? acording to LIATE

OpenStudy (turingtest):

I don't see why it makes a difference, should work both ways...

OpenStudy (turingtest):

just thinking about it... let's see

OpenStudy (turingtest):

\[\int e^{-st}\sin(kt)dt=-\frac1ke^{-st}\cos(kt)+\frac sk\int e^{-st}\cos(kt)dt\]from now on let's just call the integral \(I\) so we get\[I=\frac1ke^{-st}\sin(kt)+\frac sk\left(-\frac1ke^{-st}\cos(kt)+\frac skI\right)\]solve for \(I\)

OpenStudy (turingtest):

\[I=\frac1ke^{-st}\sin(kt)-\frac s{k^2}e^{-st}\cos(kt)+\frac {s^2}{k^2}I\]\[(1-\frac{s^2}{k^2})I=\frac1ke^{-st}\sin(kt)-\frac s{k^2}e^{-st}\cos(kt)\]\[(k^2-s^2)I=e^{-st}[k\sin(kt)-s\cos(kt)]\]\[I={e^{-st}[k\sin(kt)-s\cos(kt)]\over(k^2-s^2)}\]hmm I seem to be missing something let me check my notes...

OpenStudy (lgbasallote):

ohh i see what i did wrong now...i didnt put a k when i did my derivatives

OpenStudy (turingtest):

when you evaluate the limit in the improper integral as \(t\to0\) it should be what you have (clearly at least the \(e^{-st}\) will disappear) but I must have made a mistake try again your way

OpenStudy (lgbasallote):

yup i got the same as yours now

OpenStudy (lgbasallote):

so what's the plan how to make it look like \[\frac{s}{s^2 + k^2}\] substitute in the limits?

OpenStudy (turingtest):

I should do it again, taking each limit each time, but I have to take a break first...

OpenStudy (lgbasallote):

i kknow e^-infty is jjust 0 so i guess i'll just jump down to the good stuff \[\huge -\frac{e^0 (k\sin (0) - s\cos(0))}{s^2 + k^2}\] \[\huge \implies -\frac{-s}{s^2+k^2}\] \[\huge \implies \frac{s}{s^2 + k^2}\] nice

OpenStudy (turingtest):

yep, I somehow had s^2-k^2 in mine I must have made a sign error somewhere nice job

OpenStudy (turingtest):

if you just added a negative to turn \(s^2-k^2\) into \(s^2+k^2\) that's a mistake though you realize?

OpenStudy (lgbasallote):

prolly in the second integral

OpenStudy (lgbasallote):

nahh i did what i did earlier...i just snuck in a k

OpenStudy (turingtest):

really you should redo the problem, taking the limit every time you find v in the integration by parts steps, not waiting until the end

OpenStudy (turingtest):

ok I trust your ability to check your own work

OpenStudy (lgbasallote):

what do you mean "taking the limit everytime you find v" and "waiting until the end"

OpenStudy (lgbasallote):

well my solution was something like\[\int e^{-st} \cos kt = -\frac{e^{-st}\cos kt}{s} - \frac ks \left[-\frac{e^{-st}\sin kt}{s} + \frac ks\int e^{-st} \cos ktdt\right]\] \[\int e^{-st} \cos kt dt + \frac{k^2}{s^2} \int e^{-st} \cos kt dt = -\frac{e^{-st} \cos kt}{s} + \frac{ke^{-st}\sin kt }{s^2}\] etc.

OpenStudy (turingtest):

you should really evaluate the integral every time you take it along the way\[\int e^{-st} \cos kt dt= -\left.\frac{e^{-st}\cos kt}{s}\right|_0^\infty - \frac ks \left.\left[\left(-\frac{e^{-st}\sin kt}{s}\right)\right|_0^\infty + \frac ks\int e^{-st} \cos ktdt\right]\]this is what I mean by that...

OpenStudy (lgbasallote):

oh..but isnt it same if you just do the whole integral then simplify then evaluate integral?

OpenStudy (turingtest):

not always, though you may be able to get away with it sometimes (maybe even most if the time)

OpenStudy (turingtest):

it's one of those little details that mathematicians like Zarkon can make sound like a crime, but probably don't affect us more engineer-minded types

OpenStudy (turingtest):

evaluating you get\[I= -\lim_{n\to\infty}\left.\frac{e^{-st}\cos kt}{s}\right|_0^n - \frac ks \left.\left[\lim_{n\to\infty}\left(-\frac{e^{-st}\sin kt}{s}\right)\right|_0^n + \frac ksI\right]\]

OpenStudy (turingtest):

evaluating you get\[I= \lim_{n\to\infty}\left.\frac{e^{-st}\cos kt}{s}\right|_n^0 - \frac ks \left.\left[\lim_{n\to\infty}\left(\frac{e^{-st}\sin kt}{s}\right)\right|_n^0 + \frac ksI\right]\]

OpenStudy (turingtest):

\[I= \frac1s-\frac1s\lim_{n\to\infty}e^{-sn}\cos kn- \frac k{s^2} \left[0-\lim_{n\to\infty}\left(e^{-sn}\sin kn\right) + kI\right]\]

OpenStudy (turingtest):

are you following this still?

OpenStudy (anonymous):

Using the method I suggested before: \[ e^{ - s t}\cos(k t) =\frac 1 2 (e^{ (i k-s) t} + e^{- (i k+s)t})\\ \int_0^\infty e^{-i s t}\cos(k t) =\frac 1 2 (\int_0^\infty e^{ (i k-s) t}dt + \int_0^\infty e^{{- (i k+s)t}})dt=\\ \frac 1 2\left[ \frac{e^{ (i k-s) t}}{i k-s } \right]_{t=0}^{\infty}-\frac 1 2\left[ \frac{e^{- (i k+s)t}}{i k+s)} \right]_{t=0}^{\infty}=\\ -\frac 1 2 \frac 1 {i k-s} + \frac 1 2 \frac 1 {i k+s} =\\ \frac{s}{(k-i s) (k+i s)}= \frac s{k^2+s^2} \]

OpenStudy (turingtest):

...or continuing from my last post, notice that the limits are zero (this should be obvious if you consider that sin and cos can only vary between -1 and 1 while e^(-st) gets smaller and smaller, but can be easily proven with either de moivre or l'hospital)\[I= \frac1s-\frac1s\cancel{\lim_{n\to\infty}e^{-sn}\cos (kn)}^{\huge0}- \frac k{s^2} \left[0-\cancel{\lim_{n\to\infty}\left(e^{-sn}\sin (kn)\right)}^{\huge0} + kI\right]\]\[I=\frac1s-\frac{k^2}{s^2}I\implies(1+\frac{k^2}{s^2})I=\frac1s\implies\left(s^2+k^2\over s^2\right)I=\frac1s\]\[I=\int_0^\infty e^{-st} \cos( kt)dt={s\over s^2+k^2}=\mathcal L\{\cos (kt)\} \]